OpenStudy (arbyj):

Someone me help solve this; -10x-9=-5x+8 I have the answer but I want to make sure its correct (Please show work)

OpenStudy (trantom):

well first I would add 5x to both sides -5x-9=8 Then add 9 to both sides -5x=17 Divide both sides by -5 x=-3.4 Now check work. -10(-3.4)-9=-5(-3.4)+8 34-9=17+8 25=25 The correct answer is: x=-3.4

OpenStudy (arbyj):

Okay thank you :)

OpenStudy (trantom):

no problem

OpenStudy (trantom):

actually yes problem... its an algebra problem xD